OpenStudy (anonymous):

True or False The solution of |3x + 12| = 18 is -10 or 2.

OpenStudy (anonymous):

@Leong so for this one would it be true?

OpenStudy (anonymous):

@dan815 @Ashleyisakitty

OpenStudy (alli6106):

it is false

OpenStudy (anonymous):

@alli6106 I'm just learning this and kinda confused on it. Could you explain why to me?

OpenStudy (alli6106):

actually it is true

OpenStudy (alli6106):

first break it up into 2 equations 3x+12=18 −(3x+12)=18

OpenStudy (alli6106):

then solve the first equation

OpenStudy (alli6106):

then solve the second equation

OpenStudy (anonymous):

oh okay. thanks!

OpenStudy (alli6106):

no problem